AC Airways offers low cost air travel between Philadelphia and Atlantic City utilizing 4 planes worth a total of $12,000,000. Each plane, capable of carrying 60 passengers, makes 6 daily trips from Philadelphia to Atlantic City and 6 back from Atlantic City to Philadelphia. The price is $100 per one-way ticket. The current load factor is 65%. The annual cost of operation is $65,000,000 (labor, fuel, marketing, etc.). The company operates 365 days per year. a. Draw an ROIC tree that incorporates all of the above information. b. Compute the current ROIC. c. What is the required load factor to gain an additional 10 percentage points in ROIC? d. What is the minimum load factor at which the company breaks even
Answer #1

ROIC represents Return On Invested Capital.

In the present case Invested Capital = $12,000,000

Revenue from the sales of tickets to passengers is very much dependent on the load factor (% 0f Capacity)

At Maximum capacity utilization Revenue = 4(no. of Planes)*60(carrying capacity)*12(no. of daily trips)*365(number of operating days in a year)*100(ticket price) =$105,120,000

Therefore at current load factor of 65%, total revenue = .65*105120000 = $68,328,000

Annual cost of operations = $65,000,000

Therefore at present % of ROIC ={ (68328000-65000000)/12000000}*100 = 27.7333

Therefore required ROIC = 37.333

Revenue for required ROIC is say X, then {(X - 65000000)/12000000}*100 = 37.7333

Solving the equation we get X=$69,528,000

Therefore the load factor = (69528000/105120000)*100 = 66.14155

For additional 10% points in ROIC increase the load factor to 66.15%

Breakeven point in this case means recovery of annual cost of operations that is revenue of $65,000,000

In terms of load factor =( 65,000,000/105120000)*100 = 61.8341%